The breadth of the rectangle is one-third its length. If its perimeter is 64cm, find its area .
Answer by JulietG(1812)

Perimeter = 2L + 2W
We're given the following information:
W = 1/3L (or, L = 3W. Same thing)
P = 64
Let's put in the values:
64 = 2(3W) + 2W
Distribute:
64 = 6W + 2W
Add the Ws:
64 = 8W
Divide by 8 on each side:
8 = W
.
Now that you have the width, you know that the length is 3x that.
8 * 3 = 24 (length)
.
Let's check the info:
2 pieces for the width + 2 pieces of length = 64
8 + 8 + 24 + 24 = 64
16 + 48 does indeed = 64.
Success~!